# An insight on some properties of high order nonstandard linear multistep methods

### v1

@@ 5.1 One-dimensional logistic growth equation
$0.1$ $1.4009\cdot 10^{-1}$ $1.1611\cdot 10^{-1}$ $1.9461\cdot 10^{-1}$
@@
$0.1$ $1.4359\cdot 10^{-1}$ $9.7188\cdot 10^{-2}$ $1.1764\cdot 10^{-1}$
@@
Table 1: The errors and the corresponding orders of the different methods using different $\varphi$ functions.
Figure 1: The errors of the different methods with different $\varphi$ functions for different values of $\Delta t$.
[…]
One surprising property is the remarkable initial speed of the convergence of the third-order method, but its order decreases to the expected value of $3$ as the timestep becomes smaller. This might be the result of the effect of the function $\varphi_{8}(\Delta t)$, since this function enables higher order (see the case of SSPMS(6,4)), but this order is decreased when $\Delta t$ is smaller (and where the difference of function $\varphi_{8}(\Delta t)$ from the usual $\varphi(\Delta t)=\Delta t$ choice is less significant). This might indicate that for carefully chosen $\varphi(\Delta t)$ functions, nonstandard multistep methods might have larger orders than their standard counterparts. Another possible reason for this is the fact that we used the exact values as the initial $s$-many values of the method - if we use only a third-order approximation for these, one can only get third-order convergence (as we can see in Section 5.2).
Figure 2: The errors of the different multistep methods for different values of $\Delta t$.
NSSPMS(4,2) NSSPMS(4,3) NSSPMS(6,4)
@@
$0.05$ $1.6660\cdot 10^{-4}$ $8.2145\cdot 10^{-4}$ $1.1541\cdot 10^{-2}$
$0.05\cdot 2^{-1}$ $6.0870\cdot 10^{-5}$ $1.4527$ $5.7502\cdot 10^{-5}$ $3.8365$ $8.1974\cdot 10^{-4}$ $3.8155$
$0.05\cdot 2^{-2}$ $1.7144\cdot 10^{-5}$ $1.8280$ $4.1033\cdot 10^{-6}$ $3.8087$ $5.3510\cdot 10^{-6}$ $3.9373$
$0.05\cdot 2^{-3}$ $4.4918\cdot 10^{-6}$ $1.9324$ $3.1262\cdot 10^{-7}$ $3.7143$ $3.4099\cdot 10^{-6}$ $3.9720$
$0.05\cdot 2^{-4}$ $1.1463\cdot 10^{-6}$ $1.9703$ $2.6326\cdot 10^{-8}$ $3.5698$ $2.1515\cdot 10^{-7}$ $3.9863$
$0.05\cdot 2^{-5}$ $2.8934\cdot 10^{-7}$ $1.9862$ $2.4865\cdot 10^{-9}$ $3.4043$ $1.3511\cdot 10^{-8}$ $3.9932$
$0.05\cdot 2^{-6}$ $7.2670\cdot 10^{-8}$ $1.9933$ $2.6035\cdot 10^{-10}$ $3.2556$ $8.4697\cdot 10^{-10}$ $3.9956$
$0.05\cdot 2^{-7}$ $1.8208\cdot 10^{-8}$ $1.9967$ $2.9433\cdot 10^{-11}$ $3.1450$ $5.4143\cdot 10^{-11}$ $3.9675$
$0.05\cdot 2^{-8}$ $4.5571\cdot 10^{-9}$ $1.9984$ $3.584\cdot 10^{-12}$ $3.0376$ $5.673\cdot 10^{-12}$ $3.2545$
@@
Table 2: The errors and the corresponding orders of the different multistep methods.
Finally, we observe the way the nonstandard methods preserve the qualitative properties of the original differential equation. By Theorems 3 and 4, we know that the nonstandard multistep method preserves the boundedness property of the original solution (since the Forward Euler method has this property for $\Delta t\leq\mathcal{B}$, see Proposition 1), and condition (11) also holds. In Figure 3, we compare the results of the nonstandard methods with their standard counterparts applied with the same step sizes, initial value $\tilde{y}=3$, and function choices $\varphi_{5}$, $\varphi_{7}$, and $\varphi_{8}$, respectively. As it can be seen, for $\Delta t_{1}=0.5$ the standard methods are highly unstable: they are not bounded from below, and (11) is also violated. However, the nonstandard ones behave as expected, even for a larger timestep: their solutions are bounded from below, and the property (11) also holds with $s=4$ (in the first two cases) and with $s=6$ (in the last case). For $\Delta t_{2}=0.02$, both methods behave as expected.
Figure 3: Methods SSPMS(4,2) and NSSPMS(4,2) (upper left), SSPMS(4,3) and NSSPMS(4,3) (upper right) and SSPMS(6,4) and NSSPMS(6,4) (lower) with timesteps $\Delta t_{1}=0.5$ and $\Delta t_{2}=0.02$. For $\Delta t_{2}$, the schemes are very close to each other, therefore, only one of them is visible.
5.2 The SEIR system
The following form of the Kermack-McKendrick model [55] describes the propagation of an illness among a given constant population:
$S^{\prime}(t)$ $=-5S(t)I(t),$ (17)
$E^{\prime}(t)$ $=5S(t)I(t)-E(t),$
$I^{\prime}(t)$ $=E(t)-I(t),$
$R^{\prime}(t)$ $=I(t),$
where $S(t),E(t),I(t)$ and $R(t)$ describe the fraction of susceptible, exposed (infected but not yet infectious), ill (infected and infectious), and recovered people, respectively. It is easy to see that if $S(t_{0}),E(t_{0}),I(t_{0}),R(t_{0})\geq 0$, then $S(t),E(t),I(t),R(t)\geq 0$ holds for every $t\in[t_{0},T]$. Moreover, $S(t)+E(t)+I(t)+R(t)=\mathcal{M}$ is also true for some constant $\mathcal{M}\in\mathbb{R}^{+}$ for every $t\in[t_{0},T]$ (an interested reader might consult [56]).
Now, let us state a result about the behavior of the forward Euler method, which will be used in the choice of the function $\varphi(x)$.
Proposition 2.
Let us apply the forward Euler method to equation (17) with stepsize $\Delta t_{FE}\leq\min\left\{\dfrac{1}{\mathcal{M}},1\right\}$. Then, the forward Euler method applied to equation (17) preserves the non-negative property of the numerical solution, and $S^{n}+E^{n}+I^{n}+R^{n}=\mathcal{M}$ also holds.
The proof of a more general case can be found in [57]. Consequently, the choice $\mathcal{B}=\mathcal{C}\min\left\{\dfrac{1}{\mathcal{M}},1\right\}$ is appropriate.
A topic that has been omitted until this point is the choice of the initial values of the numerical method. In [49], it was shown that nonstandard versions of Runge-Kutta methods (i.e. ones, where the usual $\Delta t$ term is interchanged with a $\varphi(\Delta t)$ term), have (at least) the same order $p$ as their ’standard’ counterparts and also preserve the linear stability of the system if the following conditions are met:
  • Let us assume that for the ’standard’ numerical method, the linear stability and positivity are preserved if $\Delta t<\varphi^{*}$. Then the first condition is $\varphi(\Delta t)<\varphi^{*}$.
  • $\varphi(\Delta t)=\Delta t+O((\Delta t)^{p+1})$.
Note that both of these conditions are met if we choose a function $\varphi$ by the ideas discussed in Section 3.3 with $\mathcal{B}=\varphi^{*}$. Moreover, in [57] it was shown that ’standard’ Runge-Kutta methods written in the strongly stability preserving form also preserve positivity and the sum of the components if $\Delta t\leq\mathcal{C}\min\left\{\dfrac{1}{\mathcal{M}},1\right\}$ where $\mathcal{C}$ is the SSP coefficient of the Runge-Kutta method. In conclusion, for the initial $11,2,\dots,s-$ 1 values we can use a nonstandard Runge-Kutta method with a suitable order. The forms of the SSP-Runge-Kutta methods used during the experiments are listed in the Appendix.
[…]
Figure 4: The errors for different choices of function $\varphi$.

### v2

@@ 5.1.1 The non-stiff case of $c=2$
Figure 4: The sufficient bound $\mathcal{B}$ and the “real” bounds for methods NSSPMS(4,2) (upper left), NSSPMS(4,3) (upper right) and NSSPMS(6,4) (lower) for the boundedness property ($\mathcal{B}^{*}$) and the weak monotonicity ($\widetilde{\mathcal{B}}$) with $c=2$.
5.1.2 The stiff case of $c=500$
In this section, we analyze the performance of the nonstandard multistep method applied to a stiff equation. For this, we consider the previous equation (15) with $c=500$. By Proposition 1, if $\varphi(x)\leq\mathcal{C}\min\left\{\frac{1}{500},\frac{1}{\tilde{y}}\right\}$, then the method preserves the boundedness and property (11) for any choice of $\Delta t$.
We run similar tests as in the case of $c=2$ - namely, we first fix the numerical method as the fourth order NSSPMS(6,4) and observe the errors of the different choices of function $\varphi$. For this, we choose $\tilde{y}=2c=1000$ and compare the errors of the different choices at $T=\frac{1}{c}=\frac{1}{500}$ with time steps $\Delta t=2^{-k}\frac{1}{10c},\;k=0,\;1,\;\dots\;9$. The reason for the seemingly small value of $T$ is that for bigger ones, all the numerical methods are close to the exact solution (since the function is almost constant for bigger values of $t$, meaning that its change is smaller than machine precision). In Figure 5, we can see that we got a plot very similar to Figure 1, so the increase of $c$ does not affect the order of the method, but the errors are significantly bigger. The corresponding values of the errors and the orders can be seen in Table 4 in Appendix B.
Figure 5: The errors of the NSSPMS(5,4) method with different $\varphi$ functions for different values of $\Delta t$ for $c=500$.
Moreover, we also examine the order of the different multistep methods when we use $\varphi_{8}$, along with the nonstandard Runge-Kutta ones. For this, we use $\tilde{y}=2c=1000$ and compare the errors of methods NSSPMS(4,2), NSSPMS(4,3) and NSSPMS(6,4) along with Runge-Kutta ones NSSPRK(2,2), NSSPRK(3,3) and NSSPRK(10,4) with timesteps $\Delta t=2^{-k}\frac{1}{10c},\;k=0,\;1,\;\dots\;9$ at $T=\frac{1}{c}=\frac{1}{500}$. In Figure 6, it can be seen again that the methods perform as expected, even for $c=500$, but with larger errors. The corresponding values of the errors and the orders can be found in Table 5 in Appendix B. The orders are similar to the ones of the case $c=2$: the order of the method NSSPMS(4,3) is considerably larger for bigger values of $\Delta t$ when we use $\varphi_{8}$, but it is reduced when we use $\varphi_{7}$, while for the Runge-Kutta ones the order is three in both cases.
Figure 6: The errors of the different nonstandard methods for different values of $\Delta t$ with $c=500$.
As it was mentioned before, the methods preserve the desired qualitative properties even for larger timesteps, but these would result in bigger errors. In Figure 7, we compare the three nonstandard multistep methods examined before. As it can be seen, while the qualitative properties are preserved in all cases, the methods produce significant errors when we use larger timesteps. Note that in this case the “standard” multistep methods would only preserve these properties when $\Delta t\leq 2\cdot 10^{-3}$, while in the nonstandard case these are still preserved for much larger timesteps ($\Delta t_{4}$ is 100 times bigger than the bound mentioned previously). In general, one can say that the use of bigger timesteps is not efficient in neither the “standard” nor the nonstandard cases, but for the nonstandard ones, the solutions at least behave as expected, which could give us some ideas about some properties of the continuous solution.
Figure 7: The numerical solutions of the logistic equation produced by methods NSSPMS(4,2) (top), NSSPMS(4,3) (middle) and NSSPMS(6,4) (bottom) with timesteps $\Delta t_{1}=10^{-5}$, $\Delta t_{2}=3\cdot 10^{-3}$, $\Delta t_{3}=10^{-3}$ and $\Delta t_{4}=0.5$ with $c=500$. The plots on the right are zoomed-in versions of the left ones.
Lastly, we also examine how close the sufficient bound $\mathcal{B}$ is to the bound under which the methods preserve the desired properties. Like in the $c=2$ case, we consider $1000$ values of $\tilde{y}$, but in this case we use the interval $10^{-3},\;1250]$, and test the methods with $1000$ timesteps from the interval $\left[\frac{1}{500},\frac{6}{500}\right]$, and we use $T=10$. As it can be seen in Figure 8, the plots are very similar to the bounds of the case $c=2$, namely, the bound for the boundedness property can be considered sharp when $\tilde{y}\in(0,500)$, while one could also use a less strict bound for $\varphi$ in the case of $\tilde{y}>500$. Also, the weak monotonicity property (11) is preserved for $\varphi$ functions with much higher bounds.
Figure 8: The sufficient bound $\mathcal{B}$ and the “real” bounds for methods NSSPMS(4,2) (upper left), NSSPMS(4,3) (upper right) and NSSPMS(6,4) (lower) for the boundedness property ($\mathcal{B}^{*}$) and the weak monotonicity ($\widetilde{\mathcal{B}}$) with $c=500$.
5.2 The SEIR system
The following form of the Kermack-McKendrick model [59] describes the propagation of an illness among a given population with a constant influx of healthy people $\Pi\in\mathbb{R}^{+}$:
$S^{\prime}(t)$ $={\color[rgb]{0,0,0}\Pi}-5S(t)I(t),$ (16)
$E^{\prime}(t)$ $=5S(t)I(t)-E(t),$
$I^{\prime}(t)$ $=E(t)-I(t),$
$R^{\prime}(t)$ $=I(t),$
where $S(t),E(t),I(t)$ and $R(t)$ describe the number of susceptible, exposed (infected but not yet infectious), ill (infected and infectious), and recovered people, respectively. It is easy to see that if $S(t_{0}),E(t_{0}),I(t_{0}),R(t_{0})\geq 0$, then $S(t),E(t),I(t),R(t)\geq 0$ holds for every $t\in[t_{0},T]$. Moreover, $S(t)+E(t)+I(t)+R(t)={\color[rgb]{0,0,0}\Pi t+\mathcal{M}}$ is also true with constant $S(t_{0})+E(t_{0})+I(t_{0})+R(t_{0})=\mathcal{M}\in\mathbb{R}^{+}$ for every $t\in[t_{0},T]$ (an interested reader might consult [60]). From now on, we use $t_{0}=0$.
Next, let us state a result about the behavior of the forward Euler method, which will be used in the choice of the function $\varphi(x)$.
Proposition 2.
Let us apply the forward Euler method to equation (16) with stepsize $\Delta t_{FE}\leq\min\left\{{\color[rgb]{0,0,0}\dfrac{1}{5\mathcal{M}}},1\right\}$. Then, the forward Euler method applied to equation (16) preserves the non-negative property of the numerical solution, and if $\Pi=0$, then $S^{n}+E^{n}+I^{n}+R^{n}=\mathcal{M}$ also holds.
The proof of a more general case can be found in [61]. Consequently, the choice $\mathcal{B}=\mathcal{C}\min\left\{{\color[rgb]{0,0,0}\dfrac{1}{5\mathcal{M}}},1\right\}$ is appropriate.
In [61] it was shown that “standard” Runge-Kutta methods written in the strongly stability preserving form preserve positivity and the sum of the components if $\Delta t\leq\mathcal{C}\min\left\{\dfrac{1}{5\mathcal{M}},1\right\}$ where $\mathcal{C}$ is the SSP coefficient of the Runge-Kutta method. Therefore, for the initial $1,2,\dots,s-1$ values we can use a nonstandard Runge-Kutta method with a suitable order. The forms of the SSP-Runge-Kutta methods used during the experiments are listed in Appendix A.
First, the case $\Pi=0$ is considered, and then the case of nonzero influx is also observed.
5.2.1 The case $\Pi=0$
In the following, we apply the same tests to this system as we did with the previous example s in Section 5.1 with choice $\Pi=0$. First, we construct a reference solution by applying the (standard) fourth-order Runge-Kutta method with a very small timestep ($\Delta t=0.1\cdot 2^{-10}\cdot 10^{-3}$). Then, we run the nonstandard multistep method NSSPMS(6,4) with $(S^{0},E^{0},I^{0},R^{0})=(0.8,0,0.2,0)$ and with the different choices of function $\varphi$ (see Section 3.3) with timesteps $\Delta t={\color[rgb]{0,0,0}0.1/2^{k},\;k=0,1,\dots,10}$. For the values at points $t=\Delta t,\;2\Delta t,\;\dots,\;5\Delta t$, we used the nonstandard version of the fourth-order SSPRK(10,4) method with function $\varphi_{8}$. Finally, we analyze the difference between these solutions and the reference solution at $T=5$. The errors and the corresponding orders can be seen in Figure 9 and the values are listed in Table 6 in Appendix B. As we can see, the methods behave as expected, and as in the case of the first example, the first-order method with the smallest error is the one produced by $\varphi_{2}$, and the second-order one with the smallest error is $\varphi_{5}$.
Figure 9: The errors for different choices of function $\varphi$ for $\Pi=0$.
Now, we consider the convergence of the multistep methods themselves: we run the tests with the same parameters, reference solution (with timestep $\Delta t=0.05\cdot 2^{-9}\cdot 10^{-3}$), value of $T$ ($T=1$), and initial values as before, but fix the choice $\varphi_{8}$ and compare the errors of NSSPMS(4,2), NSSPMS(4,3) and NSSPMS(6,4), along with the nonstandard Runge-Kutta methods NSSPRK(2,2), NSSPRK(3,3) and NSSPRK(10,4) with timesteps $\Delta t={\color[rgb]{0,0,0}0.05/2^{k},\;k=0,1,\dots,9}$. For the initial values of the multistep methods, we used NSSPRK(2,2) with $\varphi_{5}$, NSSPRK(3,3) with $\varphi_{7}$, and NSSPRK(10,4) with $\varphi_{8}$, which are of order $2$, 33, and $4$, respectively. The errors are plotted in Figure 10 and the values of the errors and orders are listed in Table 7 in Appendix B. As it can be seen, unlike the case of the logistic equation, here both third order methods attain an order of $4$, even for smaller timesteps when we use $\varphi_{8}$ (although for the Runge-Kutta methods, this effect is less significant for smaller timesteps) - however, upon choosing $\varphi_{7}$, the order reduces to three. Moreover, the errors of the Runge-Kutta methods are smaller than the ones of the multistep schemes.
Figure 10: The errors of the different nonstandard methods for different values of $\Delta t$ for $\Pi=0$.
For the next test concerning the conservative SEIR model ($\Pi=0$), we compare the behavior of the nonstandard methods and the standard counterparts for two different timesteps (the unmentioned parameters remain the same) for $t\in[0,15]$ (in the cases of first order and third order), and $t\in[0,45]$ in the case of second order methods. For the initial values of the multistep methods, Runge-Kutta methods with an appropriate order are used: standard ones for the standard multistep methods and nonstandard ones (with appropriate choices of $\varphi$) for the nonstandard versions. We also plot the graphs produced by the nonstandard Runge-Kutta methods. In all three cases, we only plot the numerical solutions of $I^{n}$, but the others also show signs of similar behavior. In Figure 11, we can see that the standard methods behave in an unstable way: they become negative, while the nonstandard ones remain in the upper half plane for the whole process, even for larger timesteps. Also, for a small timestep, the solutions are very close to each other and all three of them behave as expected. It should also be mentioned that the errors are significantly smaller in the case of the Runge-Kutta methods, especially in the fourth order case.
Figure 11: The comparison of the standard and nonstandard methods. Upper left: SSPMS(4,2), NSSPMS(4,2) and NSSPRK(2,2) with $\Delta t_{1}=15/20$ and $\Delta t_{2}=0.02$. Upper right: SSPMS(4,3), NSSPMS(4,3) and NSSPRK(3,3) with $\Delta t_{1}=15/25$ and $\Delta t_{2}=0.02$. Lower: SSPMS(6,4), NSSPMS(6,4) NSSPRK(10,4) with $\Delta t_{1}=15/20$ and $\Delta t_{2}=0.02$. The solutions corresponding to time-step $\Delta t_{2}$ are very close to each other in all three cases.
Lastly, the sharpness of the bound for function $\varphi$ is tested in these cases too. For this, we used the initial condition $(S^{0},E^{0},I^{0},R^{0})=(1-I^{0},0,I^{0},0)$, and the methods were observed for $1000$ different values of $I^{0}$ in the interval $[0.001,0.999]$. For a given value of $I^{0}$, the method is considered to behave as expected if the given property is preserved for $1000$ different timesteps from the interval $\Delta t\in[0.5,3]$ on the time interval $t\in[0,100]$. By using the bisection method, we determine the bound $\mathcal{B}^{*}$ for the boundedness of the solution (meaning that $S^{n},I^{n},E^{n},R^{n}\in[0,1])$ and bound $\widetilde{\mathcal{B}}$ for the weak monotonicity property (11). The bounds are plotted in Figure 12. As it can be seen, considerably larger bounds for function $\varphi$ could be used - however, one might be able to find such initial conditions where the sufficient bound proved in Proposition 2 is closer to the “real” one.
Figure 12: The sufficient bound $\mathcal{B}$ and the “real” bounds for methods NSSPMS(4,2) (upper left), NSSPMS(4,3) (upper right) and NSSPMS(6,4) (lower) for the boundedness property ($\mathcal{B}^{*}$) and the weak monotonicity ($\widetilde{\mathcal{B}}$) with $\Pi=0$.
5.2.2 The case of nonzero $\Pi$
In this section, we consider the case when the SEIR model has an influx of healthy individuals, i.e. $\Pi>0$ in equation (16). It turns out that the order of method NSSPMS(6,4) upon using different $\varphi$ functions is very similar to the one we got in the case $\Pi=0$ (Figure 9 and Table 6). Moreover, the orders of the different nonstandard multistep and Runge-Kutta methods are also similar: the third order NSSPMS(4,3) and NSSPRK(3,3) (when used with $\varphi_{8}$) attain an order of four in these cases too, although numerical experiments indicate that the order of these methods decrease as $\Pi$ is increased, and gets close to $3$ when $\Pi$ is large (e.g. around $\Pi=100$).
Because of the similarities with the previous section, the main question we would like to focus on in the case of nonzero $\Pi$ is the behavior of the sum $S^{n}+E^{n}+I^{n}+R^{n}$. It is easy to see that from (16), for the function $N(t)=S(t)+E(t)+I(t)+R(t)$ (which can be thought of as the total population), we get the equation
$N^{\prime}(t)=\Pi,$
which has a solution $N(t)=\Pi t+N(0)$ (again, $t_{0}=0$). Here we observe how well the nonstandard multistep (and Runge-Kutta) methods approximate this function when applied to equation (16). In Figure 13, the sum $S^{n}+E^{n}+I^{n}+R^{n}$ is plotted for different nonstandard methods in the cases $\Pi=0$ and $\Pi=0.1$ (by using different timesteps) with initial condition $(S^{0},E^{0},I^{0},R^{0})=(0.8,0,0.2,0)$. As it can be seen, the methods preserve the sum (up to machine precision) in the case of $\Pi=0$ even with a bigger timestep ($\Delta t=1$). For the case $\Pi=0.1$, the method NSSPRK(10,4) is always close to the exact value, while the others are furher, but as we decrease the timestep, they get closer and closer to the exact value. The nonstandard Runge-Kutta methods perform better in all of the cases, while the fourth-order multistep method NSSPMS(6,4) is the furthest from the exact value.
Figure 13: The sum $S^{n}+E^{n}+I^{n}+R^{n}$ for $\Pi=0$ with $\Delta t=1$ (upper left), and for $\Pi=0.1$ with $\Delta t=0.5$ (upper right), $\Delta t=0.1$ (lower left) and $\Delta t=0.02$ (lower right) with different nonstandard methods. Method NSSPRK(10,4) is always very close to the exact solution, therefore not visible on the graphs.
6 Conclusions and discussion
In this work, we considered nonstandard multistep methods. By using a nonstandard version of the classical Taylor series (Lemma 1), we were able to show that (under some conditions for the denominator function $\varphi$) the nonstandard versions of these methods have the same order as their standard counterparts. Note that Lemma 1 might also be useful in the analysis of other, high-order nonstandard methods that are constructed in a similar fashion, i.e., involving a generalized denominator function.
In Section 3.2, it was shown that (under some further conditions for function $\varphi$) the nonstandard methods preserve the boundedness properties (and as a special case, positivity) of the original, continuous problem. To achieve this goal, a bound for the timestep should be calculated under which the Euler method preserves these properties. We note that the elementary stable property was not considered, but theorems similar to those in [51] might be proven in the future for nonstandard multistep methods too. These results would involve the analysis of the elementary stability property of standard multistep methods, which might turn out to be harder (in general, these are not even true [52], as mentioned in Section 2) - this could also be a topic for further research. On the other hand, the analysis of forward Euler methods might turn out to be easier, and for some systems, the preservation of positivity ensures elementary stability (see, for example, the case of an integro-differential equation [62] and its delayed version [63]).
In Section 5, we performed some tests for two test problems and observed that the nonstandard linear multistep methods behave as expected. In further articles, the same methods could also be applied to more robust ordinary systems and also to partial differential equations. One could also compare the performance of these multistep methods with other, higher-order methods (for example, the ones mentioned in Section 1). In the present work the nonstandard multistep methods were compared with nonstandard Runge-Kutta methods [51]. It turned out that although nonstandard multistep methods are able to achieve the same order as the nonstandard Runge-Kutta ones, the errors of the former are usually higher.
Another interesting fact was the high order of convergence in the case of the method NSSPMS(4,3). It turned out that upon applying a denominator function $\varphi$ which might be able achieve an order of four, these methods (which are usually only of third order) were also able to achieve this higher order of convergence. Because of this, it is usually advised to use such denominator functions that enable such high orders.
As it was mentioned in Section 2, from the usual two guiding principles of nonstandard methods, only one of them (the use of denominator functions) was adopted, while the implementation of nonlocal approximations was omitted. In a future work it would be interesting to incorporate the latter ideas too, which might result in smaller errors, while still keeping the high orders of the methods discussed in this paper.
Acknowledgements The author would like to thank Prof. István Faragó for his insightful comments during the preparation of the article.
Author Contributions The author is the sole contributor to this paper.
Funding: This research has been supported by the National Research, Development and Innovation Office – NKFIH, grant no. K137699. The research reported in this paper is also part of project no. BME-NVA-02, implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021 funding scheme.
Data Availability The data used to support the findings of this study are available from the corresponding author upon request.
Declarations
Conflict of interest: The authors declare no competing interests.
Appendix A. The coefficients of the different methods
  • Explicit SSPMS(4,2) [64], with $\mathcal{C}=\frac{2}{3}$:
  • Explicit SSPMS(4,3) [64], with $\mathcal{C}=\frac{1}{3}$:
  • Explicit SSPMS(6,4) [65, 66], with $\mathcal{C}=0.1648$:
    $\tilde{\alpha}_{1}=0.342460855717007,\quad\tilde{\beta}_{1}=2.078553105578060,$
    $\tilde{\alpha}_{4}=0.191798259434736,\quad\tilde{\beta}_{4}=1.164112222279710,$
    $\tilde{\alpha}_{5}=0.093562124939008,\quad\tilde{\beta}_{5}=0.567871749748709,$
    $\tilde{\alpha}_{6}=0.372178759909247.$
  • Explicit SSPRK(2,2) with $\mathcal{C}=1$ [67]:
    $u^{(1)}$ $=u^{n}+\Delta tf(u^{n}),$
  • Explicit SSPRK(3,3) with $\mathcal{C}=1$ [54]:
    $u^{(1)}$ $=u^{n}+\Delta tf(u^{n}),$
    $u^{(2)}$ $=\dfrac{3}{4}u^{n}+\dfrac{1}{4}u^{(1)}+\dfrac{1}{4}\Delta tf(u^{(1)}),$
  • Explicit SSPRK(10,4) with $\mathcal{C}=6$ [66]:
    $u^{(1)}$ $=u^{n}+\dfrac{1}{6}\Delta tf(u^{n}),$
    $u^{(j+1)}$ $=u^{(j)}+\dfrac{1}{6}\Delta tf(u^{(j)}),\qquad j=1,2,3,$
    $u^{(5)}$ $=\dfrac{3}{5}u^{n}+\dfrac{2}{5}u^{(4)}+\dfrac{1}{15}\Delta tf(u^{(4)}),$
    $u^{(j+1)}$ $=u^{(j)}+\dfrac{1}{6}\Delta tf(u^{(j)}),\qquad j=5,\dots,8,$
Appendix B. Tables of errors and orders
On the next pages, we list the errors and orders of the different methods discussed before.
@@ Appendix B. Tables of errors and orders
$0.1$ $4009\cdot 10^{-1} - 1.1$$1611\cdot 10^{-1} -$ $1.9461\cdot 10^{-1}$ -
@@
$0.1$ $1.4359\cdot 10^{-1}$ - $9.7188\cdot 10^{-2}$ - $1.1764\cdot 10^{-1}$ -
@@
Table 2: The errors and the corresponding orders of method NSSPMS(6,4) using different $\varphi$ functions while solving the logistic equation with $c=2$, plotted in Figure 1.
NSSPMS(4,2) NSSPMS(4,3) with $\varphi_{8}$ NSSPMS(4,3) with $\varphi_{7}$
@@
$0.05$ $1.6660\cdot 10^{-4}$ - $8.2145\cdot 10^{-4}$ - $3.4349\cdot 10^{-3}$ -
$0.05\cdot 2^{-1}$ $6.0870\cdot 10^{-5}$ $1.4527$ $5.7502\cdot 10^{-5}$ $3.8365$ $4.5630\cdot 10^{-4}$ $2.9122$
$0.05\cdot 2^{-2}$ $1.7144\cdot 10^{-5}$ $1.8280$ $4.1033\cdot 10^{-6}$ $3.8087$ $5.8507\cdot 10^{-5}$ $2.9633$
$0.05\cdot 2^{-3}$ $4.4918\cdot 10^{-6}$ $1.9324$ $3.1262\cdot 10^{-7}$ $3.7143$ $7.4020\cdot 10^{-6}$ $2.9826$
$0.05\cdot 2^{-4}$ $1.1463\cdot 10^{-6}$ $1.9703$ $2.6326\cdot 10^{-8}$ $3.5698$ $9.3074\cdot 10^{-7}$ $2.9915$
$0.05\cdot 2^{-5}$ $2.8934\cdot 10^{-7}$ $1.9862$ $2.4865\cdot 10^{-9}$ $3.4043$ $1.1668\cdot 10^{-7}$ $2.9958$
$0.05\cdot 2^{-6}$ $7.2670\cdot 10^{-8}$ $1.9933$ $2.6035\cdot 10^{-10}$ $3.2556$ $1.4607\cdot 10^{-8}$ $2.9979$
$0.05\cdot 2^{-7}$ $1.8208\cdot 10^{-8}$ $1.9967$ $2.9433\cdot 10^{-11}$ $3.1450$ $1.8273\cdot 10^{-9}$ $2.9990$
$0.05\cdot 2^{-8}$ $4.5571\cdot 10^{-9}$ $1.9984$ $3.5845\cdot 10^{-12}$ $3.0376$ $2.2860\cdot 10^{-10}$ $2.9988$
NSSPRK(2,2) NSSPRK(3,3) with $\varphi_{8}$ NSSPRK(3,3) with $\varphi_{7}$
$\Delta t$ errors orders errors orders errors orders
$0.05$ $3.2621\cdot 10^{-4}$ - $2.0710\cdot 10^{-6}$ - $1.4771\cdot 10^{-5}$ -
$0.05\cdot 2^{-1}$ $7.7614\cdot 10^{-5}$ $2.0714$ $2.8654\cdot 10^{-7}$ $2.8535$ $1.8598\cdot 10^{-6}$ $2.9896$
$0.05\cdot 2^{-2}$ $1.9039\cdot 10^{-5}$ $2.0274$ $3.7559\cdot 10^{-8}$ $2.9315$ $2.3330\cdot 10^{-7}$ $2.9949$
$0.05\cdot 2^{-3}$ $4.7220\cdot 10^{-6}$ $2.0114$ $4.8041\cdot 10^{-9}$ $2.9668$ $2.9213\cdot 10^{-8}$ $2.9975$
$0.05\cdot 2^{-4}$ $1.1763\cdot 10^{-6}$ $2.0051$ $6.0734\cdot 10^{-10}$ $2.9837$ $3.6548\cdot 10^{-9}$ $2.9987$
$0.05\cdot 2^{-5}$ $2.9358\cdot 10^{-7}$ $2.0024$ $7.6316\cdot 10^{-11}$ $2.9925$ $4.5709\cdot 10^{-10}$ $2.9993$
$0.05\cdot 2^{-6}$ $7.3336\cdot 10^{-8}$ $2.0012$ $9.5060\cdot 10^{-12}$ $3.0051$ $5.7211\cdot 10^{-11}$ $2.9981$
$0.05\cdot 2^{-7}$ $1.8327\cdot 10^{-8}$ $2.0006$ $1.0607\cdot 10^{-12}$ $3.1638$ $7.2802\cdot 10^{-12}$ $2.9743$
$0.05\cdot 2^{-8}$ $4.5807\cdot 10^{-9}$ $2.0003$ $1.2346\cdot 10^{-13}$ $3.1029$ $1.1662\cdot 10^{-12}$ $2.6422$
NSSPRK(6,4) NSSPRK(10,4)
$\Delta t$ errors orders errors orders
$0.05$ $1.1541\cdot 10^{-2}$ - $8.9811\cdot 10^{-9}$ -
$0.05\cdot 2^{-1}$ $8.1974\cdot 10^{-4}$ $3.8155$ $5.5896\cdot 10^{-10}$ $4.0061$
$0.05\cdot 2^{-2}$ $5.3510\cdot 10^{-6}$ $3.9373$ $3.4863\cdot 10^{-11}$ $4.0030$
$0.05\cdot 2^{-3}$ $3.4099\cdot 10^{-6}$ $3.9720$ $2.1829\cdot 10^{-12}$ $3.9974$
$0.05\cdot 2^{-4}$ $2.1515\cdot 10^{-7}$ $3.9863$ $1.4166\cdot 10^{-13}$ $3.9457$
$0.05\cdot 2^{-5}$ $1.3511\cdot 10^{-8}$ $3.9932$ $2.6867\cdot 10^{-14}$ $2.3985$
$0.05\cdot 2^{-6}$ $8.4697\cdot 10^{-10}$ $3.9956$ $4.0412\cdot 10^{-14}$ -
$0.05\cdot 2^{-7}$ $5.4143\cdot 10^{-11}$ $3.9675$ $7.4607\cdot 10^{-14}$ -
$0.05\cdot 2^{-8}$ $5.673\cdot 10^{-12}$ $3.2545$ $1.7186\cdot 10^{-13}$ -
@@
Table 3: The errors and the corresponding orders of the different multistep methods while solving the logistic equation with $c=2$, plotted in Figure 2.